OpenStudy (anonymous):

A population of rabbits is growing exponentially! In just 5 weeks it goes from a population of 100 to 1000. What is the doubling time of this population? As in, how long does the population take to double?

OpenStudy (michele_laino):

which formula do you use in order to model the exponential growth?

OpenStudy (michele_laino):

since there are more than one formula

OpenStudy (michele_laino):

we can use this formula: \[\huge N = A{e^{Bt}}\] where \(A, \;B\) are two numeric constants Now using data you provided, we have: \[\huge \left\{ \begin{gathered} 100 = A \hfill \\ 1000 = A{e^{5B}} \hfill \\ \end{gathered} \right.\] the solution of such system is: \[\huge A = 100,\quad B = 0.46\]

OpenStudy (michele_laino):

Let's suppose that at week \(\tau\), the number of rabbits, is 200, then, in order to answer to your question, we have to solve such equation: \[\huge 200 = 100 \cdot {e^{0.46 \cdot \tau }}\]